OpenStudy (anonymous):

Someone please solve this for me. solve by completing the square. 3x^2-6x=17 idk wat to do after (3x+3)^2=27 (not even sure if that is correct)

OpenStudy (anonymous):

you cannot complete the square that way.

OpenStudy (anonymous):

How come? If that is the case, would it be better to factor this problem??

OpenStudy (anonymous):

the leading coefficient has to be one, so you have to divide by 3 first

OpenStudy (cwrw238):

first take out 3 3(x^2 - 2x) = 17

OpenStudy (anonymous):

\[3x^2-6x=17\] \[x^2-2x=\frac{17}{3}\] \[(x-1)^2=\frac{17}{3}+1=\frac{20}{3}\] \[x-1=\pm\sqrt{\frac{20}{3}}\] \[x=1\pm\sqrt{\frac{20}{3}}\]

OpenStudy (cwrw238):

(x^2 - 2x) = 17/3 satellite has it

OpenStudy (anonymous):

you can factor out a 3 if you like, but you have an equation so you are allowed to divide both sides by 3 to make the leading coefficient one